# Characterizing the nuclear models informed by PREX and CREX: a view from Bayesian inference

### v1

@@ Abstract
In Fig. 3, we present the distributions of $\{\Delta F^{Ca48},\Delta F^{Pb208}\}$ colored by $b^{\prime}_{4}$ in both RMF and Skyrme models. Interestingly, we observe that an increase of $b^{\prime}_{4}$ (which represents the isovector contribution to the spin-orbit interactions) leads to a larger $\Delta F^{Pb208}$ while resulting in a smaller $\Delta F^{Ca48}$. This causes the points of ${\Delta F^{Ca48},\Delta F^{Pb208}}$ to collectively move towards the mean of joint PREX and CREX measurements. Neutron skin thickness $\Delta R_{np}$ follows the same trend as $\Delta F$ when varying $b^{\prime}_{4}$. The detailed proton and neutron density profiles are presented in the supplemental material. We can fit the parameter $b^{\prime}_{4}$ as:
$b^{\prime}_{4}=a^{\prime\prime}\Delta F^{Ca48}+b^{\prime\prime}\Delta F^{Pb208}+c^{\prime\prime},$ (14)
@@
Conclusion and discussion – In this letter, we first demonstrated that by using the existing forms of nuclear models, it’s possible to find a set of RMF parameterizations that falls within the 1-$\sigma$ of the PREX and CREX respectively, while maintaining good agreement ($<1\%$ deviation) with the binding energy, the charge radii of ${}^{208}\mathrm{Pb}$, ${}^{90}\mathrm{Zr}$ and ${}^{48}\mathrm{Ca}$ as well as tidal deformability and maximum mass of neutron stars.
@@
Based on large samples of RMF and Skyrme models from our Bayesian inference, we further identified a linear equation describing the relationship between the symmetry energy parameters ($S_{\mathrm{V}}$ and $L$) and the $\Delta F$ values of 208Pb and 48Ca. This linear relationship exhibits a weak dependency on specific choices of model parameterizations. Unlike $L$, which is positively related to the $\Delta F$ values of both the 208Pb and 48Ca, $S_{\mathrm{V}}$ impacts 208Pb and 48Ca differently. This suggests that $S_{\mathrm{V}}\gtrsim 40$ MeV is favored by PREX and CREX measurements. The large $S_{\textrm{V}}$ compared with empirical value (around 32 MeV) demonstrates the tension between the current nuclear models and joint PREX and CREX experiments.
@@
Besides $S_{\mathrm{V}}$, we found that the isovector parts of the spin-orbit interactions corresponding to the $b^{\prime}_{4}$ term in the Skyrme model also influence the neutron skin of 208Pb and 48Ca differently. We derived the equivalent density-independent $b^{\prime}_{4}$ parameter in terms of $g_{\delta}$ and $g_{\rho}$ in the RMF model and observed a similar trend. Although the correlation between $b^{\prime}_{4}$ and $\Delta F^{\mathrm{Ca}48/\mathrm{Pb}208}$ may not be strong, as also observed in Reinhard et al. (2022b), $b^{\prime}_{4}$ can be quite sensitive to a linear combination of $\Delta F^{\mathrm{Ca}48}$ and $\Delta F^{\mathrm{Pb}208}$, such as $a^{\prime\prime}\Delta F^{\mathrm{Ca}48}+b^{\prime\prime}\Delta F^{\mathrm{Pb}208}$, and can effectively control the discrepancy between the model predictions of ${\Delta F^{\mathrm{Pb208}},\Delta F^{\mathrm{Ca48}}}$ and the combined PREX and CREX measurements. We also realized that in RMF models, the increase of $b^{\prime}_{4}$ usually comes with the amplification of the (isovector) nucleon density fluctuation in the nuclei.
In RMF models, the large $b^{\prime}_{4}$ is achieved by large $g_{\delta}$ and $g_{\rho}$, which break the common quadratic dependence of symmetry energy on asymmetry. For example, the CPREX1 RMF model, calibrated to reach the $1-\sigma$ bound of PREX and CREX, ended up with $S_{\mathrm{V}}=32.9$ MeV around SNM, but $S_{\mathrm{V}}=54.3$ MeV defined with PNM. Although the slope $L$ is close to zero, the neutron star EOS in $\beta$-equilibrium remains stable due to the large L defined with PNM. CPREX1 and CPREX2 could be interesting for both the nuclear and astrophysics communities, as the neutron star properties satisfy maximum mass and tidal deformability constraints.
@@
Follow-up neutron skin experiments, like the Mainz Radius EXperiment (MREX), will continue to refine the possible uncertainty of the PREX experiment Sfienti . The linear relations between the isovector parameters ($S_{\mathrm{V}}$, $L$, and $b^{\prime}_{4}$) and the $\Delta F$ values of 208Pb and 48Ca may be extended to other double magic nuclei, serving as a simple “emulator” for rapidly and reliably approximating $S_{\mathrm{V}}$, $L$, and $b^{\prime}_{4}$ in the investigations of future neutron skin thickness experiments.
@@ Acknowledgements.
We acknowledge Jorge Piekarewicz and Jim Lattimer for the helpful discussions. T.Z. acknowledges helpful discussions held during the INT Workshop INT-22R-2A. TZ was supported by the Department of Energy, Award No. DE-FG02-93ER40756 and by the Network for Neutrinos, Nuclear Astrophysics and Symmetries, through the National Science Foundation Physics Frontier Center Grant No. PHY-2020275. M.P. acknowledges support by the Department of Energy, Award No. DE-FG02-93ER40756. ZL and AWS were supported by NSF PHY 21-16686. AWS was also supported by the Department of Energy Office of Nuclear Physics. B.K. acknowledges partial support from the Department of Science and Technology, Government of India, with grant no. CRG/2021/000101. All the Bayesian samples and data for making plots are available from the GitHub repositoryZhao .
I Supplemental Mateiral
I.1 Relativistic mean field model and Skyrme model
@@
$\Delta F$ []  0.0277$\pm$0.0055 0.0551 0.0564 0.0490 0.0435 0.0413 0.0391 0.0527 0.0330 0.0345 0.0364 0.0335 0.0362
Table 2: Experimental data for the binding energy per nucleonWang et al. (2012), charge radiiAngeli and Marinova (2013), neutron skins (excluding PREX and CREX)Lattimer (2023), charge from factor and form factor difference from PREXAdhikari et al. (2021b) for 208Pb and CREXAdhikari et al. (2022b) for 48Ca. Also displayed are the theoretical results obtained with NL3Lalazissis et al. (1997b), FSUGold2Chen and Piekarewicz (2014), IOPB-IKumar et al. (2018), IUFSUFattoyev et al. (2010), BigAppleFattoyev et al. (2020), HPUCSharma et al. (2023), BSRVKumar et al. (2023), DINOa-cReed et al. (2024) and the two new parameterizations, CPREX1 and CPREX2.
@@
Table 3: Saturation properties and neutron star properties of RMF models listed in Table 2. Saturation properties for SNM and PNM are defined in the letter. Neutron star properties are calculated with the crust EOSs constructed with the compressible liquid droptlet model repectively for various RMF models with fixed surface tension parameters $\sigma_{s}=1.2$ MeV fm-2, $S_{S}=48$ MeVLattimer and Swesty (1991).
@@
In this section, we provide an overview of the theoretical framework utilized. For a more comprehensive understanding, interested readers can refer to the detailed discussions available in various sources such as Kumar et al. (2017, 2018); Yang and Piekarewicz (2020) along with their cited references. The starting point for our relativistic calculation of nuclear response is based on the covariant model presented in Ref. Chen and Piekarewicz (2014). Additionally, we incorporate the $\delta$-meson to accommodate the CREX experiment, which prefers a softer symmetry energy Adhikari et al. (2022b). The interacting Lagrangian density can be expressed as:
@@
The Lagrangian density used in our study includes the isodoublet nucleon field $\psi$ as the basic degree of freedom. These nucleons interact via photon ($A_{\mu}$) exchange and through the exchange of four massive mesons: a scalar-isoscalar ($\sigma$), scalar-isovector ($\delta$), a vector-isoscalar ($\omega^{\mu}$), and a vector-isovector ($\rho^{\mu}$). This formulation is based on the works of Walecka (1974); Serot and Walecka (1986, 1997). The coupling constant $\Lambda_{\omega\rho}$ for the isoscalar-isovector interaction plays a crucial role in determining the density dependence of the symmetry energy, particularly its slope at saturation density $L$ as discussed in Horowitz and Piekarewicz (2001). It is noteworthy that the model can be calibrated using an analytic one-to-one correspondence between bulk parameters of infinite nuclear matter and the various coupling constants. Specifically, the values of the symmetry energy $S_{\mathrm{V}}$, its slope $L$, and curvature $K_{sym}$ at saturation density are correlated to the isovector parameters ($g_{\delta}$, $g_{\rho}$, and $\Lambda_{\omega\rho}$) as explained in Chen and Piekarewicz (2014); Singh et al. (2014). Table 4 shows the $\sigma$ meson mass and 8 coupling constants for CPREX1 and CPREX2 proposed in this letter. The masses of other massive mesons and the free nucleon are $m_{\delta}=980$ MeV, $m_{\omega}=782.5$ MeV, $m_{\rho}=763$ MeV and $m=939$ MeV. The nuclear properties of 48Ca and 208Pb are shown in Table 2 along with experimental measurement and our Hatree calculations for various models for comparison. Table 3 shows the infinite nuclear matter properties at saturation density and the neutron star properties calculated from EOSs in $\beta$-equilibrium. The crust EOSs are constructed with the compressible liquid droplet model with fixed surface tension parameters $\sigma_{s}=1.2$ MeV fm-2, $S_{S}=48$ MeVLattimer and Swesty (1991). The tabulated complete EOSs of these RMF models can be found in the GitHub repository [].
To gain a complementary understanding of both PREX and CREX measurements, we also study the Skyrme interactions inferred from those neutron skin thickness measurements. The Skyrme Hamiltonian for infinite nuclear matter is written as,
@@
To describe a finite nucleus, two additional terms are included in the Skyrme Hamiltonian Reinhard and Flocard (1995a); Chabanat et al. (1998). These terms are:
$\begin{split}H_{\mathrm{SO}}=b_{4}\textbf{J}\cdot\triangledown n+b^{\prime}_{4}(\textbf{J}_{n}\cdot\triangledown n_{n}+\textbf{J}_{p}\cdot\triangledown n_{p}),\end{split}$ (19)
@@
I.2 The approximated spin-orbit potential in RMF models
The large component of Dirac equation follows,
@@ I.2 The approximated spin-orbit potential in RMF models
where $S=g_{\sigma}\ \sigma_{0}\pm g_{\delta}\delta_{0}$, $V=g_{\omega}\ \omega\pm g_{\rho}\rho_{0}$, are mean field potential from classical expectation value of meson fields. By reducing the spherically symmetric Dirac equation with scalar and vector potentials to the non-relativistic Schrödinger equation to order $(p/m)^{2}$orowitz and Serot (1981); Reinhard (1989), the effective single nucleon classical Hamiltonian is,
@@
where $\vec{l}=\vec{r}\times\vec{p}$, $\vec{s}=\vec{\sigma}/2$ and $M^{*}=m-S(r)$ assuming that the relativistic single particle energy $\epsilon+S(r)-V(r)<<2(m-S)$. However, it’s possible to rearrange terms of higher order in $(p/m)^{2}$ to obtain an equivalent form, $M^{*}=m-V(r)$, as shown in the Appendix of Reinhard (1989). Although these two expressions are equivalent to order $(p/m)^{2}$, their ratio can be of the order of 10%. Therefore, it’s reasonable to consider the average of these two choices,
@@
which can also be justified by taking $\epsilon<<2m-S-V$ Ring (1996); Ebran et al. (2016). Since scalar and vector potentials are mainly determined by the corresponding densities of the two-form current, modified by Coulomb interaction and the self and mixing couplings of mesons, we can simplify by considering only Yukawa coupling to generate the potentials. Furthermore, vector densities are strongly correlated with scalar densities, allowing us to approximate scalar densities with vector densities:
@@
where the upper sign is for protons and the lower sign for neutrons. The difference of isospin potential between proton and neutron is,
@@
A similar form without the delta meson has been used to study the isospin dependence of the spin-orbit interaction, contrasting with the spin-orbit force in non-relativistic Skyrme models Sharma et al. (1995); Lalazissis et al. (1998); Pearson (2001). Isospin dependence of spin-orbit interaction is introduce to Skyrme models as:
@@
By comparing Eq. (34-35) with Eq. (31-32), the model parameter $b_{4}$ and $b_{4}^{\prime}$ can expressed as:
@@
In RMF models, we obtain $b_{4}$ and $b_{4}^{\prime}$ similar to those in non-relativistic models, as described in the above equations. In principle, $b_{4}$ and $b_{4}^{\prime}$ are density dependent in RMF models. We take the leading order term so that $b_{4}$ and $b_{4}^{\prime}$ are essentially a linear combination of coupling Yukawa coupling strength.
I.3 Weak Charge form factors with spin-orbit contributions
Form factors are key observable in electron scattering experiments. To calculate the weak charge form factor measured in PREX and CREX. We need to perform finite nuclei calculations to obtain the proton and neutron profiles then perform Fourier transform on the profile to get the corresponding form factor. In this work, we consider the impact of spin-orbit currents on the electromagnetic and weak charge form factors of ${}^{48}\mathrm{Ca}$ and ${}^{208}\mathrm{Pb}$. It has been shown the contributions from spin-orbit currents are comparable to the present CREX experimental error barsHorowitz and Piekarewicz (2012b), and therefore should be necessary to incorporate spin-orbit currents in the calculation of weak form factors as:
@@ I.3 Weak Charge form factors with spin-orbit contributions
where the $\mathrm{G}^{i}_{\mathrm{E,M}}$ are electric and magnetic single nucleon form factors and the $\tilde{\mathrm{G}}^{i}_{\mathrm{E,M}}$ are electric and magnetic form factors for the weak-neutral current. Similar as Horowitz and Piekarewicz (2012b), we adopted a simple dipole parametrization of these single-nucleon form factors.
@@
and
[…]
where $g_{n\kappa}(r)$ and $f_{n\kappa}$ are upper and lower component of Dirac spinor in RMF models. Wave functions in non-relativistic Skyrme model corresponds to the larger upper component $g_{n\kappa}(r)$. The lower component $f_{n\kappa}$ is obtained by assuming a free space relation Horowitz and Piekarewicz (2012b); Adhikari et al. (2022b).
$f_{n\kappa}=\frac{1}{2\mathrm{m}}(\frac{d}{dr}+\frac{\kappa}{r})g_{n\kappa}(r).$
Table 5: Prior distribution of Skyrme and RMF parameters. $n_{s}$, $B$, $M^{*}$, $K$ and $S_{V}$ map to $t_{0}$, $t_{1}$, $t_{2}$, $t_{3}$, $\gamma$ for Skyrme models and $g_{\sigma}$, $g_{\omega}$, $g_{\rho}$, $\kappa$, $\lambda$ for RMF models.
parameter prior
$n_{s}$ [MeV] [0.14,0.165]
$B$ [MeV] [-15.5,-16.5]
Both $M^{*}$ [MeV] [0.5,0.8]$\times$ 939
$K$ [MeV] [210,250]
$S_{V}$ ([MeV] [20,50]

### v2

@@ Abstract
In Fig. 3, we present the distributions of $\{\Delta F^{\mathrm{Ca}48},\Delta F^{\mathrm{Pb}208}\}$ colored by $b^{\prime}_{4}$ in RMF and Skyrme models. As $b^{\prime}_{4}$ increases, $\Delta F^{\mathrm{Pb}208}$ rises while $\Delta F^{\mathrm{Ca}48}$ drops, moving predictions closer to the PREX-2/CREX mean. Neutron skin thickness $\Delta R_{np}$ follows the same trend. The detailed proton and neutron density profiles are presented in the supplemental material. We can fit the parameter $b^{\prime}_{4}$ as:
$b^{\prime}_{4}=a^{\prime\prime}\Delta F^{\mathrm{Ca}48}+b^{\prime\prime}\Delta F^{\mathrm{Pb}208}+c^{\prime\prime},$ (12)
@@
Conclusion and discussion – In this letter, we systematically investigated the influence of $S_{\mathrm{V}}$, L and $b^{\prime}_{4}$ on the $\Delta F$ of $^{48}\mathrm{Ca}$ and $^{208}\mathrm{Pb}$ based on Bayesian inference.
@@
We identified a quasi linear relationship between the symmetry energy parameters ($S_{\mathrm{V}}$ and $L$) and the $\Delta F$. This relationship exhibits a weak dependency on nuclear models. Unlike $L$, which is positively related to the $\Delta F$ values of both the 208Pb and 48Ca, $S_{\mathrm{V}}$ impacts 208Pb and 48Ca differently. This suggests that $S_{\mathrm{V}}\gtrsim 40$ MeV is favored by PREX-2/CREX measurements. The large $S_{\textrm{V}}$ compared with empirical value (around 32 MeV) suggests the tension between the current nuclear models and joint PREX-2/CREX experiments.
@@
Additionally, we found that the isovector parts of the spin-orbit interactions in both Skyrme and RMF models influence the neutron skin of 208Pb and 48Ca differently. Recently, the authors of Ref. Yue et al. (2024) observed similar behaviors in the extended Skyrme models. Although the correlation between $b^{\prime}_{4}$ and $\Delta F^{\mathrm{Ca}48/\mathrm{Pb}208}$ may not be strong, as observed in Reinhard et al. (2022b), $b^{\prime}_{4}$ can be quite sensitive to a linear combination of $\Delta F^{\mathrm{Ca}48}$ and $\Delta F^{\mathrm{Pb}208}$, such as $a^{\prime\prime}\Delta F^{\mathrm{Ca}48}+b^{\prime\prime}\Delta F^{\mathrm{Pb}208}$, and by tuning $b^{\prime}_{4}$ the agreement between theoretical predictions and joint PREX-2/CREX measurements is efficiently improved. However, We realized that in RMF models, the increase of $b^{\prime}_{4}$ usually comes with the amplification of the (isovector) nucleon density fluctuation, altering of single particle energy spectrum as discussed in Kunjipurayil et al. (2025), and an increase of $g_{\delta}$ and $g_{\rho}$ which break the common quadratic dependence of symmetry energy on asymmetry. The Skyrme / RMF posterior dataset may help in the search of optimized models with more physical constraints in future investigations.
Finally, among numerous RMF models that fall within the 1-$\sigma$ of the PREX-2/CREX respectively, we identify CPREX1 that maintains good agreement ($1\%$ deviation) with the binding energy, the charge radii of $^{208}\mathrm{Pb}$, $^{90}\mathrm{Zr}$ and $^{48}\mathrm{Ca}$ as well as tidal deformability and maximum mass of NSs. Similar to DINO with large $g_{\delta}$ and $g_{\rho}$, CPREX1 deviates from the quadratic dependence of symmetry energy on isospin asymmetry, yielding $S_{\mathrm{V}}=32.9$ MeV around SNM, but $S_{\mathrm{V}}=54.3$ MeV when defined with respect to PNM.
@@
Follow-up neutron skin experiments, like the Mainz Radius EXperiment (MREX), will continue to refine the possible uncertainty of the PREX-2 experiment Sfienti . The linear relations between the isovector parameters ($S_{\mathrm{V}}$, $L$, and $b^{\prime}_{4}$) and the $\Delta F$ values of 208Pb and 48Ca may serve as a simple “emulator” for rapidly and reliably approximating $S_{\mathrm{V}}$, $L$, and $b^{\prime}_{4}$ in the investigations of neutron skin thickness experiments.
@@ Acknowledgements.
We acknowledge Jorge Piekarewicz and Jim Lattimer for the helpful discussions during the INT Workshop INT-22R-2A. T.Z. was supported by the Network for Neutrinos, Nuclear Astrophysics and Symmetries, through the National Science Foundation Physics Frontier Center Grant No. PHY-2020275. M.P. and T.Z acknowledges support by the Department of Energy, Award No. DE-FG02-93ER40756. Z.L. and A.W.S. were supported by NSF PHY 21-16686. A.W.S. was also supported by the Department of Energy Office of Nuclear Physics. B.K. acknowledges partial support from the Department of Science and Technology, Government of India, with grant no. CRG/2021/000101. All the Bayesian samples and data for making plots are available from the GitHub repositoryZhao . TZ and ZL contributed equally to this work.
Supplemental Material
.1 Relativistic mean field model and Skyrme model
@@
[…]
Table 2: Experimental data for the binding energy per nucleonWang et al. (2012), charge radiiAngeli and Marinova (2013), neutron skins (excluding PREX-2/CREX)Lattimer (2023), charge from factor and form factor difference from PREXAdhikari et al. (2021b) for 208Pb and CREXAdhikari et al. (2022b) for 48Ca. Also displayed are the theoretical results obtained with NL3Lalazissis et al. (1997b), FSUGold2Chen and Piekarewicz (2014), IOPB-IKumar et al. (2018), IUFSUFattoyev et al. (2010), BigAppleFattoyev et al. (2020), HPUCSharma et al. (2023), BSRVKumar et al. (2023), DINOa-cReed et al. (2024) and the two new parameterizations, CPREX1 and CPREX2.
@@
Table 3: Saturation properties and neutron star properties of RMF models listed in Table 2. Saturation properties for SNM and PNM are defined in the letter. For all models except DINOa–c and CPREX1–2, the two definitions coincide. In these cases, the discrepancy signals the breakdown of the quadratic approximation of symmetry energy. Neutron star properties are calculated with the crust EOSs constructed with the compressible liquid droptlet model repectively for various RMF models with fixed surface tension parameters $\sigma_{s}=1.2$ MeV fm-2, $S_{S}=48$ MeVLattimer and Swesty (1991).
@@
In this section, we provide an overview of the theoretical framework utilized. For a more comprehensive understanding, interested readers can refer to the detailed discussions available in various sources such as Kumar et al. (2017, 2018); Yang and Piekarewicz (2020) along with their cited references. The starting point for our relativistic calculation of nuclear response is based on the covariant model presented in Ref. Chen and Piekarewicz (2014). Additionally, we incorporate the $\delta$-meson to accommodate the CREX experiment, which prefers softer symmetry energy Adhikari et al. (2022b). The interacting Lagrangian density can be expressed as:
@@
The Lagrangian density used in our study includes the isodoublet nucleon field $\psi$ as the basic degree of freedom. These nucleons interact via photon ($A_{\mu}$) exchange and through the exchange of four massive mesons: a scalar-isoscalar ($\sigma$), scalar-isovector ($\delta$), a vector-isoscalar ($\omega^{\mu}$), and a vector-isovector ($\rho^{\mu}$). This formulation is based on the works of Walecka (1974); Serot and Walecka (1986, 1997). The coupling constant $\Lambda_{\omega\rho}$ for the isoscalar-isovector interaction plays a crucial role in determining the density dependence of the symmetry energy, particularly its slope at saturation density $L$ as discussed in Horowitz and Piekarewicz (2001). It is noteworthy that the model can be calibrated using an analytic one-to-one correspondence between bulk parameters of infinite nuclear matter and the various coupling constants. Specifically, the values of the symmetry energy $S_{\mathrm{V}}$, its slope $L$, and curvature $K_{sym}$ at saturation density are correlated to the isovector parameters ($g_{\delta}$, $g_{\rho}$, and $\Lambda_{\omega\rho}$) as explained in Chen and Piekarewicz (2014); Singh et al. (2014). Table 4 shows the $\sigma$ meson mass and 8 coupling constants for CPREX1 and CPREX2 proposed in this letter. The masses of other massive mesons and the free nucleon are $m_{\delta}=980$ MeV, $m_{\omega}=782.5$ MeV, $m_{\rho}=763$ MeV, and $m=939$ MeV. The nuclear properties of 48Ca and 208Pb are shown in Table 2 along with experimental measurement and our Hartree calculations for various models for comparison. The nuclear properties of 48Ca and 208Pb are shown in Table 2, along with experimental measurements and our Hartree calculations for various models for comparison. These two nuclei, together with 90Zr (not shown), are selected for inclusion in our Bayesian likelihood to limit computational cost, as the evaluation of finite nuclei properties constitutes the main bottleneck in the sampling. While the CPREX models reproduce the binding energies and charge radii of these nuclei with high accuracy, they are not optimized for global nuclear fits. For example, the binding energy per nucleon of 16O for CPREX1 is predicted with an error of about 2%, which, while larger than for the nuclei included in the fit, remains competitive with or better than the uncertainty range of chiral EFT (1–10%). We emphasize that prioritizing consistency with PREX2/CREX and astrophysical constraints necessarily limits the model’s ability to fit additional nuclei, given its finite degrees of freedom. Fitting to a broader range of nuclei would risk overfitting and reduce the sensitivity to the observables we aim to constrain. Table 3 shows the infinite nuclear matter properties at saturation density and the neutron star properties calculated from EOSs in $\beta$-equilibrium. The crust EOSs are constructed with the compressible liquid droplet model with fixed surface tension parameters $\sigma_{s}=1.2$ MeV fm-2, $S_{S}=48$ MeVLattimer and Swesty (1991). The tabulated complete EOSs of these RMF models can be found in the GitHub repositoryZhao .
To gain a complementary understanding of both PREX-2/CREX measurements, we also study the Skyrme interactions inferred from those neutron skin thickness measurements. The Skyrme Hamiltonian for infinite nuclear matter is written as,
@@
To describe a finite nucleus, two additional terms are included in the Skyrme Hamiltonian Reinhard and Flocard (1995a); Chabanat et al. (1998a). These terms are:
$\begin{split}H_{\mathrm{SO}}=b_{4}\textbf{J}\cdot\nabla n+b^{\prime}_{4}(\textbf{J}_{n}\cdot\nabla n_{n}+\textbf{J}_{p}\cdot\nabla n_{p}),\end{split}$ (17)
@@
.2 The approximated spin-orbit potential in RMF models
The large component of the Dirac equation follows,
@@
where $S=g_{\sigma}\ \sigma_{0}\pm g_{\delta}\delta_{0}$, $V=g_{\omega}\ \omega\pm g_{\rho}\rho_{0}$, are mean-field potential from classical expectation value of meson fields. By reducing the spherically symmetric Dirac equation with scalar and vector potentials to the non-relativistic Schrödinger equation to order $(p/m)^{2}$Horowitz and Serot (1981); Reinhard (1989), the effective single nucleon classical Hamiltonian is,
@@
where $\vec{l}=\vec{r}\times\vec{p}$, $\vec{s}=\vec{\sigma}/2$ and $M^{*}$ could be defined as the Dirac mass $M^{D}=m-S(r)$ assuming that the relativistic single particle energy $\epsilon+S(r)-V(r)<<2(m-S)$. However, it’s possible to rearrange terms of higher order in $(p/m)^{2}$ to obtain an equivalent form, $M^{*}=m-V(r)$, as shown in the Appendix of Reinhard (1989). Although these two expressions are equivalent to order $(p/m)^{2}$, their ratio can be of the order of 10%, as $S(0)-V(0)\approx 50$ MeV, the empirical depth of the nucleon potential well. Therefore, we consider the average of these two choices originally introduced in Koepf and Ring (1991),
@@
which can also be justified by taking $\epsilon<<2m-S-V$ Ring (1996); Ebran et al. (2016). Note that this definition of $M^{*}$ should not be confused with the Dirac or Landau effective mass. Our qualitative conclusions remain unchanged regardless of the specific definition of $M^{*}$ used. Since scalar and vector potentials are mainly determined by the corresponding densities of the two-form current, modified by Coulomb interaction and the self and mixing couplings of mesons, we can simplify by considering only Yukawa coupling to generate the potentials. Furthermore, vector densities are strongly correlated with scalar densities, allowing us to approximate scalar densities with vector densities:
@@
where the upper sign is for protons and the lower sign is for neutrons. The difference of isospin potential between proton and neutron is,
@@
A similar form without the delta meson has been used to study the isospin dependence of the spin-orbit interaction, contrasting with the spin-orbit force in non-relativistic Skyrme models Sharma et al. (1995); Lalazissis et al. (1998); Pearson (2001). Isospin dependence of spin-orbit interaction is introduced to Skyrme models as:
@@
By comparing Eq. (32-33) with Eq. (29-30), the model parameter $b_{4}$ and $b_{4}^{\prime}$ can expressed as:
@@
In RMF models, we obtain $b_{4}$ and $b_{4}^{\prime}$ similar to those in non-relativistic models, as described in the above equations. In principle, $b_{4}$ and $b_{4}^{\prime}$ are density-dependent in RMF models. We take the leading order term so that $b_{4}$ and $b_{4}^{\prime}$ are essentially a linear combination of coupling Yukawa coupling strength.
.3 Weak Charge form factors with spin-orbit contributions
Form factors are key observable in electron scattering experiments. To calculate the weak charge form factor measured in PREX-2/CREX. We need to perform finite nuclei calculations to obtain the proton and neutron profiles then perform Fourier transform on the profile to get the corresponding form factor. In this work, we consider the impact of spin-orbit currents on the electromagnetic and weak charge form factors of $^{48}\mathrm{Ca}$ and $^{208}\mathrm{Pb}$. It has been shown the contributions from spin-orbit currents are comparable to the present CREX experimental error barsHorowitz and Piekarewicz (2012b), and therefore should be necessary to incorporate spin-orbit currents in the calculation of weak form factors as:
@@
where $\tau=q^{2}/4m^{2}$, the $\mathrm{G}^{i}_{\mathrm{E,M}}$ are electric and magnetic single nucleon form factors and the $\tilde{\mathrm{G}}^{i}_{\mathrm{E,M}}$ are electric and magnetic form factors for the weak-neutral current. Similar to Horowitz and Piekarewicz (2012b), we adopted a simple dipole parametrization of these single-nucleon form factors.
@@
$\mathrm{F}_{\mathrm{V}}(q)=\sum_{n\kappa}(2j+1)\int_{0}^{\infty}[g^{2}_{n\kappa}(r)+f^{2}_{n\kappa}(r)]j_{0}(qr)dr~,$
and
$\mathrm{F}_{\mathrm{T}}(q)=\sum_{n\kappa}2(2j+1)\int_{0}^{\infty}g_{n\kappa}(r)f_{n\kappa}(r)j_{1}(qr)dr~,$
where $g_{n\kappa}(r)$ and $f_{n\kappa}(r)$ are the upper and lower components of Dirac spinor in RMF models. Wave functions in the non-relativistic Skyrme model correspond to the larger upper component $g_{n\kappa}(r)$. The lower component $f_{n\kappa}(r)$ is obtained by assuming a free space relation Horowitz and Piekarewicz (2012b); Adhikari et al. (2022b).
$f_{n\kappa}(r)=\frac{1}{2\mathrm{m}}(\frac{d}{dr}+\frac{\kappa}{r})g_{n\kappa}(r).$
Table 5: Prior distribution of Skyrme and RMF parameters. $n_{s}$, $B$, $m^{*}$ (or $M^{D}$), $K$ and $S_{V}$ map to $t_{0}$, $t_{1}$, $t_{2}$, $t_{3}$, $\gamma$ for Skyrme models and $g_{\sigma}$, $g_{\omega}$, $g_{\rho}$, $\kappa$, $\lambda$ for RMF models.
parameter prior
$n_{s}$ [MeV] [0.14,0.165]Drischler et al. (2024)
$B$ [MeV] [-15.5,-16.5]Drischler et al. (2024)
@@
$K$ [MeV] [210,250] Xu et al. (2021)
$S_{V}$ ([MeV] [20,50] Sun et al. (2024)